OpenStudy (anonymous):

Walker is buying a house for $225,000. He is financing $195,000 and obtained a 30-year, fixed-rate mortgage with a 6.625% interest rate. How much are his monthly payments?

OpenStudy (anonymous):

So first of all you need to find out what the 6.625% of 195.000 is. 195.000*.06625=12918.75 Then add it on 195.000 to get the total amount to pay for. 195.000+12918,75=207.918,75 Then find out how many months you have to pay it off and then divide it by the total payment. 30*12(months)=360months 207.918,75/360=577.55$ (his monthly payments)